OpenStudy (anonymous):

Select the system of linear equations that has the solution (3, 0).

OpenStudy (cpinder):

Any answer choices?

OpenStudy (anonymous):

These are multiple choice answers

OpenStudy (anonymous):

1. X+y=7 2. 3x+y=9 x-3y=3 3.x+3y=3 X+y=3 4. X-y=7 x+7y=3 x+y=5

OpenStudy (anonymous):

Please help me

OpenStudy (cpinder):

What is your x and what is your y with the ordered pair?

Directrix (directrix):

Look at the options: #1 shows only 1 equation and should show two. #5 shows 3 equations and should show two. @Cruznatalie150

OpenStudy (unklerhaukus):

check the remaining options by substituting and x= 3, y=0 If the left hand side of each equation equals the right, then that equation has a solution (3,0) You need all equations in the system to have the solution (3,0).